OpenStudy (matimaticas):

Trig question When doing Polar Form of a Complex Number, if the point plotted is in quadrant 2 or 3, do you add pi?

OpenStudy (matimaticas):

@jim_thompson5910 please help

jimthompson5910 (jim_thompson5910):

what is the point?

OpenStudy (matimaticas):

4/-3

jimthompson5910 (jim_thompson5910):

(4, -3) ?

OpenStudy (matimaticas):

yeah its in the form z = a + bi (-3 + 4i)

jimthompson5910 (jim_thompson5910):

(4,-3) = (a,b) so a+bi = 4-3i

jimthompson5910 (jim_thompson5910):

assuming (4,-3) is in the form (a,b) where a+bi

OpenStudy (matimaticas):

my bad. In this case it's -3 + 4i

jimthompson5910 (jim_thompson5910):

is the initial point (4,-3) or (-3,4)

OpenStudy (matimaticas):

(-3,4)

jimthompson5910 (jim_thompson5910):

and you want to convert this to polar form right?

OpenStudy (matimaticas):

yes

OpenStudy (jdoe0001):

one may note and angle in quadrant II woudln't amount to a \(\pi\) unles is \(\pi\) itself

jimthompson5910 (jim_thompson5910):

(x,y) = (-3,4) x = -3 y = 4 now use the formula \[\Large r = \sqrt{x^2+y^2}\] to find the value of r

OpenStudy (matimaticas):

so r = 5

jimthompson5910 (jim_thompson5910):

yes r = 5

jimthompson5910 (jim_thompson5910):

to find theta, you'll use \[\Large \theta = \arctan\left(\frac{y}{x}\right)\] you need to add on pi (I think this is what you were initially asking about) because you will start off in Q4, but adding pi gets you to Q2 where the angle is

jimthompson5910 (jim_thompson5910):

so again, the arctan function in this case will return an angle in Q4 adding pi to that result brings the angle to Q2

OpenStudy (matimaticas):

YES that is what I was confused about, when you add pi to theta

jimthompson5910 (jim_thompson5910):

so hopefully that clears things up?

OpenStudy (matimaticas):

so you have to add pi when its in Q1 or Q4?

jimthompson5910 (jim_thompson5910):

The range of arctan is -pi/2 < y < pi/2 so that means that the outputs of arctan are either in Q1 or Q4

jimthompson5910 (jim_thompson5910):

if you get something in Q4, but the angle is really in Q2, you have to add pi but if the angle is really in Q4, then you just leave it as it is

jimthompson5910 (jim_thompson5910):

same for Q1 and Q3

OpenStudy (matimaticas):

ok.. thank you :)

jimthompson5910 (jim_thompson5910):

np